INFRARED SPECTRA OF CARBONYL SULFIDE-ACETYLENE TRIMERS: OCS-(C_{2}H_{2})_{2} AND TWO ISO-MERS OF (OCS)_{2}-C_{2}H_{2}

\underline{MAHIN AFSHARI}, M. DEHGHANY, J. N. OLIAEE, N. MOAZZEN-AHMADI, Department of Physics and Astronomy, University of Calgary, Calgary, AB T2N 1N4, Canada; A.R.W. MCKEL-LAR, Steacie Institute for Molecular Sciences, National Research Council of Canada, Ottawa, ON K1A 0R6, Canada.

Spectra of acetylene-carbonyl sulphide trimers in the region of the OCS ν_{1} fundamental (∼2062 cm^{−1}) are observed using a tunable diode laser to probe a pulsed supersonic slit jet expansion. A previous microwave study of (OCS)_{2}- C_{2}H_{2} by Peebles and Kuczkowski^{a} gave a nonplanar triangular twisted structure, which could be thought of as a polar OCS dimer plus a C_{2}H_{2} monomer lying above the dimer plane. In the present work, three infrared bands are analyzed. The first band clearly belongs to this previously known (OCS)_{2}-C_{2}H_{2} complex. The second band can be assigned as an isomer of (OCS)_{2}-C_{2}H_{2} having a similar structure, but with a nonpolar OCS dimer plus a C_{2}H_{2} monomer above the dimer plane. The third band is assigned to OCS-(C_{2}H_{2})_{2}. The rotational constants and dipole moment components of all three bands are consistent with barrel shape structures having C_{1}, C_{2} and C_{s} symmetries, respectively.

aS.A. Peebles and R.L. Kuczkowski, J. Chem. Phys. 111, 10511 (1999).